The mean salary paid to 500 employees working in a firm was found to be Rs. 180.40. After disbursement of salaries for a certain month, it was discovered that the salary of two employees was wrongly entered as Rs. 297 and Rs. 165 against their correct salary of Rs. 197 and Rs. 185, respectively. Find the correct mean salary.
1
Expert's answer
2017-02-20T13:49:11-0500
Let's denote the sum of all employee salaries as S, therefore we have:
S/500 = 180.4

Let's denote the correct mean salary as x, therefore equation for x:
x = (S - 297 + 197 - 165 + 185) / 500 = S/500 - 80/500 = 180.4 - 0.16 = 180.24

The correct mean salary is equal to 180.24